OpenStudy (anonymous):

I'm going to try this one more time, please help as my brain is completely fried, I am trying to solve a simple equation for x .............. x/3 - 7/8 = 2/5(x + 15/4) What I have so far is this: x/3 - 7/8 = 2/5(x + 15/4) x/3 - 7/8 = 2/5x + 3/2 x/3 - 7/8 + 7/8 = 2/5x + 3/2 + 7/8 x/3 = 2/5x + 51/40 Now here is where I got a little stuck, x/3 - 2/5x = 2/5x - 2/5x + 51/40 ? = 51/40 Am I wrong so far, this is how we did it for practice?

OpenStudy (amistre64):

why do you insist on beating your head against fractions?

OpenStudy (amistre64):

clear them out and ply with the integers

OpenStudy (anonymous):

I don't understand

OpenStudy (amistre64):

you dont understand what ???

OpenStudy (amistre64):

fractions are intrinsically difficult to work with simply because we do not "think" in fractions ....

OpenStudy (amistre64):

recall that an equation remains balanced by doing the same thing to both sides ....

OpenStudy (anonymous):

Let me try to explain, I am 43 years old, I have not taken an algebra class since 9th grade and I barely made it through that. I am a graphic designer working on getting my BA who has absolutely NO USE for any of this, but I have to take it. This teacher is big on formulas, and on the practice problems we did on Tuesday, this was the method we were using for all the other problems so I assumed it was the same for this, I DON'T UNDERSTAND THIS STUFF, so when you ask me questions, I'm sitting here with a blank expression on my face........

OpenStudy (amistre64):

i turned 40 last week, and im going to college to get a BA in Math :)

OpenStudy (anonymous):

I'm with you on the I don't think in fractions, I SUCK at fractions, but I don't know how else to do this

OpenStudy (anonymous):

YIKES, YOUR INSANE

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

thats besides the point lol

OpenStudy (anonymous):

I honestly don't have a clue what I am doing, but it's already 12:19 am my time, and I have to get up for work at 6 am, this is killing me

OpenStudy (anonymous):

and I'm only on question 2 of 10

OpenStudy (anonymous):

for the record, this is now the second time I am trying to pass this class, I didn't even come close the first time

OpenStudy (amistre64):

x/3 - 7/8 = 2/5(x + 15/4) ; lets be smart about this and clear out the fractions x/3 - 7/8 = 2/5(x + 15/4) ; lets start by multiplying both sides by 3 x.3/3 - 7.3/8 = 2.3/5(x + 15/4) ; since 3/3 = 1 that "clears" the fraction x - 21/8 = 6/5(x + 15/4) ; and we keep doing that; this time by "8" 8.x - 21.8/8 = 6.8/5(x + 15/4) ; 8/8 = 1 and clears the fraction 8x - 21 = 48/5(x + 15/4) ; and again, but by 5 this time 8.5x - 21.5 = 48.5/5(x + 15/4) ; since 5/5 = 1 .... 40x - 105 = 48(x + 15/4) ; now we do it one last time to clear the "4" 40.4x - 105.4 = 48.4(x + 15/4) 160x - 420 = 192(4x + 15) now we are out of fraction

OpenStudy (amistre64):

the numbers are big, but id take big over fractured any day

OpenStudy (amistre64):

160x - 420 = 768x + 2880 -2880 -2880 -------------------------- 160x-2300 = 768x -160x -160x ------------------- -2300 = 608x x = -2300/608

OpenStudy (amistre64):

since they are even we can split them in half over and over again :) 2300/2 = 1150/2 = 575 ------ ------- ---- 608/2 304/2 152 so if we are lucky, x = -575/152 but we would have to dbl chk to make sure

OpenStudy (amistre64):

x = -285/8 in the end

OpenStudy (anonymous):

good grief, the double check is going to kill me....lol

OpenStudy (amistre64):

i see my mistake, its towards the end; I actually used that "4" twice ...

OpenStudy (amistre64):

40.4x - 105.4 = 48.4(x + 15/4) 160x -420 = 48(4x+15) 160x -420 = 192x+720 -720 -720 ----------------------- 160x-1140 = 192x -160x -160x ------------------- -1140 = 32x x = -1140/32 = -570/16 = -285/8

OpenStudy (amistre64):

now that works out :)

OpenStudy (anonymous):

ty, I'm going to try to write all that out the correct way.............good luck to me...lol

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

thanks, I'm going to need it, I have to pass this class eventually, I just hope I can do it this time because it's the last one

OpenStudy (amistre64):

your time is on the east coast

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

im in florida

OpenStudy (anonymous):

me too...lol Ft. Myers

OpenStudy (amistre64):

Tampa ish

OpenStudy (anonymous):

OMG, you have to come down here and help me.....lol

OpenStudy (amistre64):

with the price of gas at 4$ lol

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

ever been to Dade City?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Great work @amistre64 and @lilangl :)

OpenStudy (amistre64):

well ive been to FtMyers back when I was able to tent houses

OpenStudy (amistre64):

uhoh, if you stand real still he gets confused and goes away :)

OpenStudy (anonymous):

not on my end....lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:-/ Where'd you go?

OpenStudy (amistre64):

im going to bed, got diffyQs and other stuff in the am ;) Gnite

OpenStudy (anonymous):

I have a feeling my test is going to be late, there is no way I'm going to get this done, it took me almost 12 hours to do 10 practice problems, I have 10 test questions but they are of course even harder........ It's just not humanly possible with a full time job and family to get something like this done in one night.

OpenStudy (anonymous):

gn ty for your help